# Contrasting exchange-field and spin-transfer torque driving mechanisms in all-electric electron spin resonance

### v1

@@ I Introduction
Coherent manipulation of individual electron spins localized in single atoms or molecules on a surface using electron spin resonance (ESR) implemented in a scanning tunneling microscope (STM) is a novel platform for the study of many-body physics with possible applications for quantum simulations using artificially constructed interacting spin systems on surfaces  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Such a system can be modeled as a magnetic impurity connected to two leads, which provide a harmonic alternating current (AC) drive and a static external magnetic field that Zeeman-splits the up and down spin states. When one of the leads is ferromagnetic, a polarized current is detected in the direct current (DC) signal. In this situation, adjusting the driving frequency of the AC voltage to correspond with the Zeeman splitting leads to resonance behavior in the electric current.
@@
The mechanism mediating the coupling between the AC electric field and the spin remains a topic of debate [26, 27, 28, 29, 30, 31]. Recent studies have proposed that the driving mechanism originates from the magnetic field component of the electric field, which is sufficiently amplified in the electromagnetic cavity formed between the tip and the magnetic impurity [30]. This driving is necessary because a time-dependent Kondo scattering Hamiltonian lacks the charge fluctuations required to provide any driving to explain the experimentally observed Rabi rates, as we mention in [27]. The presence of an AC magnetic field as a product of the AC electric field contrasts with recent experimental findings [21, 22, 20], which suggest that a magnetic atom must be placed near the spin impurity to enable driving outside the tunnel junction but still under the influence of the electric field generated by the tip body. A long-range electromagnetic driving field appears inconsistent with the requirement for a localized magnetic field source.
@@
More recently, a Kondo Hamiltonian approach was used to derive an effective magnetic field proportional to the tip’s spin polarization and the AC voltage, which can drive the magnetic impurity [31]. This model accounts for current flow via multi-electron (cotunneling) processes in a region where sequential tunneling is suppressed by Coulomb blockade [32, 33, 27]. While this proposed driving mechanism aligns with the approach used in this work, it does not fully capture all features of the experimentally observed homodyne detection. Homodyne detection arises from the misalignment between the STM tip’s spin and the magnetic impurity and the AC driving, imposing a change on the tunneling magnetoresistance, as observed in ESR measurements. In particular, Homodyne detection predicts the strongest ESR signal when the tip spin polarization and the magnetic impurity are mostly parallel [34, 21], whereas [31] differs from previous reports by showing the maximum ESR signal at $90^{\circ}$.
Some of us have proposed an all-electric model for ESR in the STM, which can efficiently drive the spin of an Anderson impurity model based on the principle of barrier modulation [35, 36]. We were able to solve the resulting quantum master equation either in the long-time limit or explore real-time spin dynamics by solving the time propagation of the Hamiltonian [37]. At present, the model is limited to the weak coupling regime, imposing sequential tunneling processes, which implies no cotunneling effects, although the model is extended to the Coulomb Blockade regime via the tail of the resonance [36]. It is also important to note that much of the existing literature on cotunneling in transport overlooks the exchange field contribution from the sequential regime. As pointed out by Braun et al. in Refs. [38, 39], this approach is inaccurate. A comprehensive theory that fully integrates the main elements of both the sequential and cotunneling regimes for a periodically driven system is still lacking, but recent works are aiming towards this goal [40, 41, 42]. Nevertheless, our extended sequential model qualitatively describes many of the experimentally observed features of ESR of single spin impurities on surfaces.
@@
However, a fundamental question was left unanswered: how does the barrier modulation couple to the spin, and can this mechanism be linked to any physical measurements? In this work, we aim to explore this question by establishing connections with related models for quantum dots connected to magnetic leads [43, 44, 45, 38, 42] where an exchange field mechanism between the tip and sample is demonstrated and applied to experimental studies [46, 47]. This comparison is justified, as atomic or molecular spin impurities can be seen as miniaturized versions of lithographically defined larger quantum dots. We will show that the ESR mechanism is highly dependent on the applied DC bias relative to the ionization and charging energy thresholds of the impurity i.e., it depends if we are above the energy thresholds or inside the Coulomb Blockade regime. Specifically, we will contrast two driving mechanisms:
@@
  • Field-like Torque which dominates ESR for DC biases in the Coulomb blockade regime. It is mediated by the exchange magnetic field induced by the spin-polarized tip, leading to a conventional ESR mechanism and enabling coherent spin manipulation via electric pulses. We refer to this driving mechanism as FLT.
@@
  • Spin-transfer Torque which drives ESR for DC biases outside the Coulomb blockade regime. It is dominated by spin accumulation on the impurity, driven by the current which torques the impurity spin and leads to strong spin polarization of the spin impurity. We refer to this driving mechanism as STT.
We emphasize that the driving mechanism is a distinct concept and exists separately from the transport regime, which is determined by the DC bias value relative to the ionization energies. The driving mechanisms (FLT, STT) are general and exist below and above the ionization thresholds. The transport regime dictates how strongly FLT and STT affect the magnetic impurity.
@@ II.1 Model Hamiltonian
The model used here is a single orbital Anderson impurity model [48] (SAIM) of a quantum impurity (QI), $H_{\mathrm{QI}}$, connected to two ferromagnetic or normal electrodes (tip or substrate) that can be treated as fermionic baths by time-dependent hoppings $H_{T}(t)$ as introduced in previous works[35, 36], see Fig. 1a. This model is general enough to describe a wide range of physical systems, where the impurity spin can be localized in point defects, atoms or molecules. The full system is given by $H(t)=H_{\rm L}+H_{\rm R}+H_{\rm QI}+H_{\rm T}(t),$ where the first two terms describes the spin polarized or normal fermionic baths, the second term is the QI and the third term contains the hopping connecting baths and QI. The bath itself can be expressed using the common form (Eq. 1), where each state is described by an electrode label $\alpha=L,R$, spin projection along the $z$ axis $\sigma=\uparrow,\downarrow$, and momentum $k$. For both ferromagnetic electrodes, we consider that their magnetizations point in the $z$ direction. In this case, the effective spin asymmetry, $D_{\alpha\uparrow}\neq D_{\alpha\downarrow}$, in the spin-dependent density of states $D_{\alpha\sigma}$ at the Fermi level can be described by the spin polarization $P_{\alpha}\equiv(D_{\alpha\uparrow}-D_{\alpha\downarrow})/(D_{\alpha\uparrow}+D_{\alpha\downarrow})$, where $D_{\alpha\sigma}=D_{\alpha}\left(1/2+\sigma P_{\alpha}\right)$ with $P_{\alpha}\in[-1,1]$ and the extreme cases representing a fully spin polarized lead pointing down, $-1$, or up, $+1$.
For the sake of simplicity, we focus on the simplest QI in Eq. 2 which contains the Coulomb repulsion $U$, the onsite energy $\varepsilon$ and an external magnetic field $\mathbf{B}=B(\sin\phi\sin\theta,\cos\phi\sin\theta,\cos\theta)$ expressed in spherical coordinates pointing in any direction, noncollinear with the direction of ferromagnetic lead magnetizations. Thus [^1] Heisenberg exchange coupling $J$ between transport spin $\mathbf{\hat{s}}$ and another localized spin or between localized spins can be seamlessly incorporated into the model [35], enabling the study of arbitrary spin clusters and more complex spin structures. Future work will explore this in depth., we have
@@
$H_{T}(t)$ $=$ $\sum_{\alpha k\sigma}\left(t_{\alpha}(t)c^{\dagger}_{\alpha k\sigma}d_{\sigma}+t_{\alpha}^{*}(t)d^{\dagger}_{\sigma}c_{\alpha k\sigma}\right).$ (3)
Each electrode is characterized by a chemical potential $\mu_{\alpha}$ and a temperature $T_{\alpha}$. Our convention for the applied bias is such that $eV_{DC}=\mu_{L}-\mu_{R}$, meaning the DC bias is applied to the left electrode which we will assign as the tip, see Fig. 1a and b. Consequently, electrons move from sample to tip at positive applied bias, with the SOMO energy being associated with $\varepsilon$ and the SUMO energy with $\varepsilon+U$, where $\varepsilon$ and $U$ are the ionization and Coulomb repulsion energies respectively. Changing the sign convention to one where the DC bias is applied to the sample, $eV_{DC}=\mu_{R}-\mu_{L}$, only requires swapping the labels for the $\varepsilon$ and $\varepsilon+U$ energies to keep the correct physical meaning of the orbitals. The right chemical potential is assumed to be nearly aligned with the Fermi energy and grounded, which goes in line with the asymmetry tunneling conditions $t_{L}>t_{R}$ in [49] and follows the experimental evidence of the impurity being more strongly coupled to the substrate than the STM tip. In this model, only the hopping terms include a periodic modulation in the form:
@@
with $A_{\alpha}\propto V_{\mathrm{AC}}$. This functional dependence has already been analyzed in previous works [27, 35, 36], so we defer a more detailed examination to a future publication. It is important to highlight that this approach does not explicitly include nor exclude the piezoelectric effect [50]; rather, it accounts for all possible physical contributions on the driving parameter.
@@
The use of a reduced density matrix formalism allows us to define any observable straightforwardly. To do so, we consider all possible charge configurations: zero, one, and two electrons in the QI (see Fig. 1b). These configurations are denoted as $|p\rangle$, with $p=g,e,0,2$. The first two are the singly occupied states of the QI at $\varepsilon$ energy, corresponding to the ground and excited states, respectively, and their time evolution fully describes the spin state of the system. The states 00 (empty) and $2$ (doubly occupied) are transient states necessary to facilitate transport through the impurity. The first state represents an empty quantum dot, which is aligned with the Fermi energy and therefore chosen to be at zero energy, $E_{0}=0$. Consequently, removing one electron from the impurity requires an energy cost of $E_{1}-E_{0}=\varepsilon$. Meanwhile, the doubly occupied state is formed when two electrons with opposite spins occupy the same level or orbital, with an energy given by $E_{2}=2\varepsilon+U$. To create this state, an electron from the leads must overcome an energy barrier of $E_{2}-E_{1}=\varepsilon+U>0$ to be placed at the singly occupied state. From the energy perspective, the quantum impurity hosts two electrons, one at energy $\varepsilon$ and another at $\varepsilon+U$, with the latter being a effective unoccupied singly state. Thus, the relevant energies for transport are the differences between states with different numbers of electrons, i.e., $\varepsilon$ and $\varepsilon+U$, as shown in Fig. 1b, rather than the absolute energy of the SIAM levels.
Since the impurity Hamiltonian does not depend on time, it is convenient to diagonalize the QI Hamiltonian and introduce the reduced density matrix. Thus, we can write $H_{\rm QI}|l\rangle=\sum_{j}E_{j}\hat{\rho}_{jj}|l\rangle=E_{l}|l\rangle.$ The Hubbard operators $\hat{\rho}_{lj}=|l\rangle\langle j|$, with the Latin characters ($l,j$ and later $u,v,\dots$) referring to eigenstates of the impurity, are closely related to the reduce density matrix [35, 36, 51]. The specific case of the eigenstates $g$ and $e$ to be $\downarrow$ and $\uparrow$, respectively, in Fig. 1b corresponds to a positive out-of-plane magnetic field, i.e., along the positive $z$-axis. As we will later mention, the external magnetic field forces the spin to be aligned parallel to it for an isotropic $g$-factor. This implies that the spin will be quantized in the magnetic field direction rather than $z$, and so the reduced density matrix, by definition, is represented in the coordinate system, associated with the magnetic field direction, where the spin is quantized and the QI Hamiltonian is diagonal. Thus, for example, the expectation value of the spin components needs to be computed with care, see Appendix A, since we cannot simply use the Pauli matrices.
The Hubbard operators will emerge in the hopping Hamiltonian as well as [52],
@@
The matrix element $\lambda_{lj,\sigma}=\langle l|d_{\sigma}|j\rangle=\langle l|0\rangle\langle\sigma|j\rangle+\langle l|\bar{\sigma}\rangle\langle 2|j\rangle$ reflects the change of the many-body configurations on the QI, they are computed in Appendix B.
II.2 Quantum master equation
At this stage, we derive a QME for the reduced density matrix by treating the coupling between the QI and the reservoirs up to the lowest order in perturbation theory in $H_{T}$ in the Born-Markov approximation [53, 54], meaning $\hbar^{-1}\Gamma\tau_{B}\ll 1$ where $\Gamma$ is proportional to the rate of change of the reduced density operator due to the coupling to the bath and $\tau_{B}$ is the characteristic decay time of the bath [55]. Following Refs. [56, 57, 58, 59, 60, 61, 51, 35, 36], one arrives at:
$\hbar\dot{\rho}_{lj}(t)$displaystyle-- $i\Delta_{lj}\rho_{lj}(t)=\sum_{vu}\left\{\left[\Gamma_{vl,ju}(t)+\Gamma^{*}_{uj,lv}(t)\right]\rho_{vu}(t)\right.$ (5)
-- $\left.\Gamma_{jv,vu}(t)\rho_{lu}(t)-\Gamma^{*}_{lv,vu}(t)\rho_{uj}(t)\right\},$
where $\Delta_{lj}=E_{l}-E_{j}$. The QME, Eq. 5, is physically meaningful in the limit of weak coupling between QI and electrodes, and, as we mentioned, it is represented in the quantization axis dictated by the spin orientation. In previous works [35], we derived a Floquet version of the QME, which has proven useful in the long-time limit for efficiently computing spectra by transforming the equations to the frequency domain. This formulation incorporates Floquet components of the density matrix, labeled by the Floquet number $n$. The rates in Eq. 5 can be written as the sum of two contributions per electrode $\alpha$:

### v2

@@ I Introduction
Coherent manipulation of individual electron spins localized in single atoms or molecules on a surface using electron spin resonance (ESR) implemented in a scanning tunneling microscope (STM) is a novel platform for the study of many-body physics with potential applications for quantum simulations using artificially constructed interacting spin systems on surfaces  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Such a system can be modeled as a magnetic impurity connected to two leads, which provide a harmonic alternating current (AC) drive and a static external magnetic field that Zeeman-splits the up and down spin states. When one of the leads is ferromagnetic, a polarized current is detected in the direct current (DC) signal. In this situation, adjusting the driving frequency of the AC voltage to correspond with the Zeeman splitting leads to resonance behavior in the electric current.
However, the mechanism mediating the coupling between the AC electric field and the spin remains a topic of debate [31, 32, 33, 34, 35, 36], as selection rules indicate that such an interaction should be forbidden. To reconcile this, some of us have proposed an all-electric model for ESR in the STM, which can efficiently drive the spin of an Anderson impurity model based on the principle of barrier modulation [37, 38]. We solved the resulting quantum master equation either in the long-time limit or by simulating real-time spin dynamics via direct time propagation of the Hamiltonian [39], enabling the implementation of quantum gates and highly entangled states [40]. At present, the model is limited to the weak coupling regime, imposing sequential tunneling processes, which implies no cotunneling effects, although the model is extended to the Coulomb Blockade regime via the tail of the resonance [38].
@@
A different line of work has proposed a time-dependent Kondo-like Hamiltonian in which the spin impurity is subjected to a driving that originates not from the Kondo coupling itself, but from the magnetic field component of the electric field, sufficiently amplified in the electromagnetic cavity formed between the tip and the spin [35]. This model, which represents a valuable contribution, requires the inclusion of this additional driving because a purely time-dependent Kondo scattering Hamiltonian lacks the necessary charge fluctuations to account for the experimentally observed Rabi rates, as discussed in [32]. In other words, it discards the exchange field contribution described in [41], which is essential in this context. However, the main concern is that the presence of an AC magnetic field as a product of the AC electric field contrasts with recent experimental findings [22, 23, 21], suggesting that a magnetic atom must be placed near the spin impurity to enable driving outside the tunnel junction but still under the influence of the electric field generated by the tip body. A long-range electromagnetic driving field appears inconsistent with the requirement for a localized magnetic field source.
@@
A subsequent approach, also based on a Kondo Hamiltonian, derived an effective magnetic field proportional to the spin polarization of the tip and the AC voltage, which can drive the magnetic impurity [36]. While this proposed driving mechanism aligns to some extent with our approach, it does not fully capture all features of the experimentally observed homodyne detection. Homodyne detection arises from the misalignment between the STM tip’s spin and the magnetic impurity under AC driving. In particular, it predicts the strongest ESR signal when the spin polarization of the tip and the magnetic impurity are mostly parallel [42, 22], whereas Ref. [36] differs from previous reports by showing the maximum ESR signal at $90$ degrees. This discrepancy is likely again related to the lack of a proper incorporation of the exchange field, an issue that we believe is also present in Ref. [31].
As has been pointed out, much of the existing literature on cotunneling (Kondo) in transport overlooks the exchange field contribution from the first-order correction in perturbation theory [43, 44, 31, 35, 36]. Indeed, the Schrieffer–Wolff transformation applied to the Anderson impurity model to obtain a higher-order cotunneling Hamiltonian neglects all first-order coupling terms, including the exchange field contribution. This was shown to be inaccurate when considering ferromagnetic electrodes, spin polarization, [45, 41] since the exchange field survives in the Coulomb blockade regime [45, 41, 46, 47]. A comprehensive theory that consistently incorporates both sequential and cotunneling processes, together with first-order coherent contributions such as the exchange field, for periodically driven systems is still lacking, although recent works have made progress toward this goal [48, 49, 46].
@@
Coming back to our theoretical proposal, a fundamental question was left unanswered: how does the barrier modulation couple to the spin, and can this mechanism be linked to any physical measurements? Does it capture the exchange field contribution? In this work, we aim to explore these questions by establishing connections with related models for quantum dots connected to ferromagnetic leads [50, 51, 52, 41, 46] where an exchange field mechanism between the tip and sample is demonstrated and applied to experimental studies [53, 54]. This comparison is justified, as atomic or molecular spin impurities can be seen as miniaturized versions of larger lithographically defined quantum dots. We will show that the ESR mechanism is highly dependent on the applied DC bias relative to the ionization and charging energy thresholds of the impurity i.e., it depends if we are above the energy thresholds or within the Coulomb Blockade regime. Specifically, we will contrast two driving mechanisms:
@@
  • Field-like Torque dominates ESR for DC biases in the Coulomb blockade regime. It is mediated by the exchange magnetic field induced by the spin-polarized tip, leading to a conventional ESR mechanism, Homodyne detection and enabling coherent spin manipulation via electric pulses. We refer to this driving mechanism as FLT.
@@
  • Spin-transfer Torque drives ESR for DC biases outside the Coulomb blockade regime. It is dominated by spin accumulation on the impurity, driven by the current, which torques the impurity spin and leads to strong spin polarization of the impurity. We refer to this driving mechanism as STT.
We emphasize that the driving mechanism is a distinct concept and exists separately from the transport regime. The driving mechanisms (FLT, STT) are general and exist below and above the ionization thresholds. The transport regime dictates how strongly FLT and STT affect the magnetic impurity. It is important to mention that both driving regimes have been recently observed experimentally in a pentacene molecule [26], and they are expected to emerge in other molecular systems with well-defined orbital levels. The current work is inspired by the combination of key physical concepts introduced in Ref. [41] and the recent experimental realizations in the ESR-STM in Ref. [26].
@@ II.1 Model Hamiltonian
The model used here is a single orbital Anderson impurity model [55] (SAIM) of a quantum impurity (QI), $H_{\mathrm{QI}}$, connected to two ferromagnetic or normal electrodes (tip or substrate) that can be treated as fermionic baths by time-dependent hoppings $H_{T}(t)$ as introduced in previous works [37, 38], see Fig. 1a. This model is general enough to describe a wide range of physical systems, where the impurity spin can be localized in point defects, atoms, or molecules. The full system is given by $H(t)=H_{\rm L}+H_{\rm R}+H_{\rm QI}+H_{\rm T}(t),$ where the first two terms describes the spin-polarized or normal fermionic baths, the second term is the QI and the third term contains the hopping connecting baths and QI:
@@
$H_{T}(t)$ $=$ $\sum_{\alpha k\sigma}\left(t_{\alpha}(t)c^{\dagger}_{\alpha k\sigma}d_{\sigma}+t_{\alpha}^{*}(t)d^{\dagger}_{\sigma}c_{\alpha k\sigma}\right),$ (3)
The bath itself can be expressed using the common form (Eq. (1)), where each state is described by an electrode label $\alpha=L,R$, spin projection along the $z$ axis $\sigma=\uparrow,\downarrow$, and momentum $k$. For both ferromagnetic electrodes, we consider that their magnetizations point in the $z$ direction. In this case, the effective spin asymmetry, $D_{\alpha\uparrow}\neq D_{\alpha\downarrow}$, in the spin-dependent density of states $D_{\alpha\sigma}$ at the Fermi level can be described by the spin polarization $P_{\alpha}\equiv(D_{\alpha\uparrow}-D_{\alpha\downarrow})/(D_{\alpha\uparrow}+D_{\alpha\downarrow})$, where $D_{\alpha\sigma}=D_{\alpha}\left(1/2+\sigma P_{\alpha}\right)$ with $P_{\alpha}\in[-1,1]$ and the extreme cases representing a fully spin polarized lead pointing down, 1-1, or up, +1+1. Each electrode is characterized by a electrochemical potential $\mu_{\alpha}$ and a temperature $T_{\alpha}$.
For the sake of simplicity, we focus on the simplest [^1] More complex QIs with Heisenberg exchange coupling $J$ between the transport spin $\mathbf{\hat{s}}$ and another localized spin, or between multiple localized spins, can be seamlessly incorporated into the model [37], enabling the study of arbitrary spin clusters and more intricate spin structures. Future work will explore this in depth. QI in Eq. (2), which includes the Coulomb repulsion $U$, the onsite energy $\varepsilon$, and an external magnetic field $\mathbf{B}=B(\sin\phi\sin\theta,\cos\phi\sin\theta,\cos\theta)$ expressed in spherical coordinates, pointing in an arbitrary direction, generally noncollinear with the ferromagnetic lead magnetization. The transport spin is defined as $\mathbf{\hat{s}}=\hbar\sum_{\sigma\sigma^{\prime}}d_{\sigma}^{\dagger}\bm{\sigma}_{\sigma\sigma^{\prime}}d_{\sigma^{\prime}}/2$, with $\bm{\sigma}$ the vector of Pauli matrices. This definition incorporates the factor of $\hbar$ from the Bohr magneton into the spin operator.
Our convention for the applied bias is such that $eV_{DC}=\mu_{L}-\mu_{R}$, meaning the DC bias is applied to the left electrode, which we will assign as the tip, see Fig. 1a and b. Consequently, electrons move from the sample to the tip at a positive applied bias. The onsite energy (ionization energy) $\varepsilon$ is associated with the singly occupied molecular orbital (SOMO), while $\varepsilon+U$ is linked to the singly unoccupied molecular orbital (SUMO). Changing the sign convention to one where the DC bias is applied to the sample, $eV_{DC}=\mu_{R}-\mu_{L}$, only requires swapping the labels for $\varepsilon$ and $\varepsilon+U$ energies to keep the correct physical meaning of the orbitals. The right electrochemical potential is assumed to be nearly aligned with the Fermi energy and grounded, which goes in line with the asymmetry tunneling conditions $t_{L}<t_{R}$ in [56] and follows the experimental evidence of the impurity being more strongly coupled to the substrate than the STM tip.
In this model, only the hopping terms, Eq. (3), include a periodic modulation in the form:
@@
with $A_{\alpha}\propto V_{\mathrm{AC}}$. This functional is justified under appropriate parametric conditions, and its dependence has already been discussed in previous works [32, 37, 38, 57]; therefore, we defer a more detailed examination to a future publication. It is important to highlight that this approach does not explicitly include nor exclude the piezoelectric effect [58]; rather, it accounts for all possible physical contributions to the driving parameter.
@@
The use of a reduced density matrix formalism allows us to define any observable straightforwardly. To do so, we consider all possible charge configurations: zero, one, and two electrons in the QI (see Fig. 1b). These configurations are denoted as $|p\rangle$, with $p=g,e,0,2$. The first two are the singly occupied states of the QI at $\varepsilon$ energy, corresponding to the ground and excited states, respectively, and their time evolution fully describes the spin state of the system. The states 0 (empty) and $2$ (doubly occupied) are transient states necessary to facilitate transport through the impurity. The first state represents an empty quantum dot, aligned with the grounded electrochemical potential $\mu_{R}$ for convenience and therefore chosen to be at zero energy, $E_{0}=0$. Consequently, removing one electron from the impurity requires an energy cost of $E_{1}-E_{0}=\varepsilon$. Meanwhile, the doubly occupied state is formed when two electrons with opposite spins occupy the same level or orbital, with an energy given by $E_{2}=2\varepsilon+U$. To create this state, an electron from the leads must overcome an energy barrier of $E_{2}-E_{1}=\varepsilon+U>0$ to be placed at the singly occupied state. From the energy perspective, the quantum impurity hosts two electrons, one at energy $\varepsilon$ and another at $\varepsilon+U$, with the latter being an effective unoccupied singly state. Thus, the relevant energies for transport are the differences between states with different numbers of electrons, i.e., $\varepsilon$ and $\varepsilon+U$, as shown in Fig. 1b, rather than the absolute energy of the SIAM levels.
Since the impurity Hamiltonian does not depend on time, it is convenient to diagonalize the QI Hamiltonian and introduce the reduced density matrix. Thus, we can write $H_{\rm QI}|l\rangle=\sum_{j}E_{j}\hat{\rho}_{jj}|l\rangle=E_{l}|l\rangle.$ The Hubbard operators $\hat{\rho}_{lj}=|l\rangle\langle j|$, with the Latin characters ($l,j$ and later $u,v,\dots$) referring to eigenstates of the impurity, are closely related to the reduce density matrix [37, 38, 59, 60]. These operators describe transitions between many‑body eigenstates $|l\rangle$ and $|j\rangle$ of the isolated quantum impurity, and reduce to a projector on a single state when $l=j$. If we assemble all these operators $\rho_{lj}=\mathrm{Tr}[\hat{\rho}_{T}(t)\hat{\rho}_{lj}]$, with $\hat{\rho}_{T}(t)$ being the density matrix of the full system given by the total Hamiltonian $H(t)$, then the coefficients $\rho_{lj}$ are exactly the elements of the impurity’s reduced density matrix in that eigenbasis. For the isolated impurity, subspaces with different electron numbers are decoupled, so $\rho_{lj}=0$ whenever $|l\rangle$ and $|j\rangle$ differ by one electron. When we couple the impurity to the electrodes, these previously disconnected subspaces become linked by single‑electron tunneling, and these operators mediate the transition probability, as the transformation of the tunneling Hamiltonian in terms of Hubbard operators indicates:
@@
The matrix element $\lambda_{lj,\sigma}=\langle l|d_{\sigma}|j\rangle=\langle l|0\rangle\langle\sigma|j\rangle+\langle l|\bar{\sigma}\rangle\langle 2|j\rangle$ reflects the change of the many-body configurations on the QI, they are computed in Appendix B. Unlike ordinary fermionic creation and annihilation operators, Hubbard operators do not obey Wick’s theorem but instead a graded Lie algebra. As a result, perturbation theory is more challenging, making, for example, a cotunneling expansion with exchange‑field corrections more cumbersome.
The specific case in Fig. 1b, where the ground and excited eigenstates, $g$ and $e$, are labeled $\downarrow$ and $\uparrow$, respectively, corresponds to a positive out-of-plane magnetic field (i.e. oriented along the positive $z$ axis). For an isotropic $g$-factor and when the coupling to the electrodes is weak, the Zeeman interaction aligns a spin-$1/2$ parallel to the magnetic-field direction. Therefore, if the external field is arbitrarily oriented the spin aligns with $\mathbf{B}$ rather than with the fixed $z$ axis, and in general $H_{\mathrm{QI}}$ is no longer diagonal in the $z$-projection basis. Accordingly, by our definition, we represent the reduced density matrix in the coordinate system where $H_{\mathrm{QI}}$ is diagonal (the QI eigenbasis), which here coincides with the magnetic-field direction. In systems with strong anisotropy or for higher-spin cases, however, the spin(s) will not in general point along $\mathbf{B}$ and the eigenbasis of $H_{\mathrm{QI}}$ need not coincide with the magnetic-field axis. Thus one must work with two coordinate systems: the fixed Cartesian frame $(x,y,z)$ in which the $z$ axis defines the computational basis (chosen parallel to the tip magnetization for convenience), and the QI eigenbasis (the axis where $H_{\mathrm{QI}}$ is diagonal). Quantities defined in the QI eigenbasis must be projected onto the fixed Cartesian frame when comparing to measurements, because the transport current probes the $z$ component of the spin since the tip magnetization is along $z$, as we will later. Consequently, expectation values of spin components must be computed with care, see Appendix A; one cannot simply insert Pauli matrices in the computational-$z$ basis without the appropriate basis rotation or change in the basis set.
II.2 Quantum master equation (QME)
At this stage, we derive a QME for the reduced density matrix by treating the coupling between the QI and the reservoirs up to the lowest order in perturbation theory in $H_{T}$ in the Born-Markov approximation [61, 62], meaning $\hbar^{-1}\Gamma\tau_{B}\ll 1$ where $\Gamma$ is proportional to the rate of change of the reduced density operator due to the coupling to the bath and $\tau_{B}$ is the characteristic decay time of the bath [63]. Following Refs. [64, 65, 66, 67, 68, 69, 59, 37, 38], we write:
$\hbar\dot{\rho}_{lj}(t)$ - $i\Delta_{lj}\rho_{lj}(t)=\sum_{vu}\left[\Gamma_{vl,ju}(t)+\Gamma^{*}_{uj,lv}(t)\right]\rho_{vu}(t)$ (5)
- $\sum_{vu}\left\{\Gamma_{jv,vu}(t)\rho_{lu}(t)+\Gamma^{*}_{lv,vu}(t)\rho_{uj}(t)\right\},$
where $\Delta_{lj}=E_{l}-E_{j}$ and every Latin character refers to eigenstates of the QI. The QME, Eq. (5), is physically meaningful in the limit of weak coupling between QI and electrodes, and, as we mentioned, it is represented in the eigenstates basis set, parallel to the external magnetic field for an isotropic spin-1/2. In a previous work [37], we derived a Floquet version of the QME, which has proven useful in the long-time limit for efficiently computing spectra by transforming the equations to the frequency domain. This formulation incorporates Floquet components of the density matrix, labeled by the Floquet number $n$. The rates [^2] While rates are typically expressed in units of $1/\text{s}$, in our formulation they are instead defined in units of energy. This reflects their interpretation as an energy width rather than a conventional rate. in Eq. (5) can be written as the sum of two contributions per electrode $\alpha$: